40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 16 01:46:41 2014, Nichin, Update, Electronics, Test run of PDFR system PDFR_testRun_15-07-2014.pdf
    Reply  Wed Jul 16 02:22:10 2014, Koji, Update, Electronics, Test run of PDFR system 
Message ID: 10212     Entry time: Wed Jul 16 01:46:41 2014     Reply to this: 10214
Author: Nichin 
Type: Update 
Category: Electronics 
Subject: Test run of PDFR system 

A test run was conducted on the PDFR system last afternoon and transimpedance plots were generated for 6 of the PDs. The laser was shut down after the test run.

I have not verified (yet) if the transimpedance values indicated by the plots are correct or not. The values mostly look INCORRECT. But the peaks are exactly where they need to be. *phew!*

Reasons: Incorrect calibration, Light other than from the PDFR system fibers on the PDs

Will have to work on debugging all this.

Attachment 1: PDFR_testRun_15-07-2014.pdf  192 kB  Uploaded Thu Jul 17 13:02:58 2014  | Hide | Hide all
PDFR_testRun_15-07-2014.pdf PDFR_testRun_15-07-2014.pdf PDFR_testRun_15-07-2014.pdf PDFR_testRun_15-07-2014.pdf PDFR_testRun_15-07-2014.pdf PDFR_testRun_15-07-2014.pdf
ELOG V3.1.3-